OpenStudy (anonymous):

Given 5n² + 6n + 7 = n² - 4n : a. Find the value of the discriminant. b. State the number and type of roots/solutions/zeros.[2 real, 2 imaginary, 1 real]

OpenStudy (nory):

First: Get all the terms on one side, and a 0 on the other.

OpenStudy (anonymous):

4n^2+10n+7=0

OpenStudy (anonymous):

@Nory

OpenStudy (nory):

Good. Now remember the discriminant formula: \[\sqrt{b ^{2}-4ac}\]

OpenStudy (nory):

Oops...I didn't need the square root there.

OpenStudy (nory):

Anyway, if the bit under the square root is positive, then it has two real solutions. If it is 0, then it has one real solution, and if it is negative, no real solutions.

OpenStudy (nory):

So figure out what a, b, and c are and plug them into the formula.

OpenStudy (anonymous):

would i use the quadratic formula

OpenStudy (nory):

That's the second step, the step where you actually _find_ the solutions. See, do you see the expression I wrote? The one with the square root? Does that look familiar? It's actually part of the quadratic formula.

OpenStudy (anonymous):

4n^2+10n+7=0 10^2-4(10)(7) 100-40(7) 100-280 -180

OpenStudy (anonymous):

its that about right?

OpenStudy (nory):

Just about. You just missed one step. When you do b^ - 4ac, the second term should be 4(4)(7) instead of 4(10)(7). But I see you have the basic idea.

OpenStudy (anonymous):

so this has 2 solutions ? and the value is 12?

OpenStudy (nory):

Well, the discriminant would be 10^2 - 4(4)(7) 100 - 16*7 And (you can use a calculator): 100 - 112 -12

OpenStudy (nory):

The discriminant is negative, so what does that tell you about the solutions?

OpenStudy (anonymous):

theres no real solution

OpenStudy (nory):

exactly. Good job. So therefore, there are 2 imaginary solutions. And you're done!

OpenStudy (nory):

good work. :)